# No-hair and almost-no-hair results for static axisymmetric black holes and ultracompact objects in astrophysical environments

### v1

@@ V Almost-no-hair results for horizonless ultracompact objects
As a final case, let us examine the most relevant situation from the astrophysical point of view. It corresponds to ultracompact stellar-like objects that can mimic most properties of general relativistic black holes, allowing them to go unnoticed as black holes. Take the innermost equipotential surface outside the matter content of the object and a sphere in Weyl coordinates $r^{2}+z^{2}=\text{constant}$ encompassing this surface. Assuming that outside this sphere we are in vacuum, the $U$ function can always be expanded in terms of Weyl multipoles originated inside the sphere.
Now, let us consider that the multipolar structure is such of a Schwarzschild black hole plus some additional contributions:
$U(r>R_{e})=U_{S}+\sum_{\ell}U^{(\ell)}.$ (58)
This multipolar expansion is not correct/valid when approaching the minimum equipotential surface with $U=\epsilon$, in regions with $r\to 0$ (notice that the equipotentials are not cylinders and hence do not correspond to $r=\text{constant}$). Then, trying to see what happens with ultracompact object as $\epsilon$ tends to 00 entails a subtle problem. There is not a single limit for how matter behave inside these surfaces and two separated families of solutions arise with well differentiated behaviours in the $\epsilon=0$ situation: the Schwarzschild family (which has a regular horizon) and the Curzon family (which displays a directional curvature singularity where $U\rightarrow-\infty$). It all depends on whether by making $\epsilon$ smaller and smaller one unveils the $1/r^{\ell+1}$ behaviour of the $U^{\ell}$ (the $\ell=0$ monopolar case for instance; see Fig. 10 for a pictorial representation of this limit) or the $\log(r)$ behaviour is enforced by making the contributions from the multipole terms to be parametrically small enough for a fixed $\epsilon$ (Schwarzschild case; see Fig. 11 for a pictorial representation). In the former case, the addition of a Schwarzschild contribution to a Curzon-like expansion can be regarded as a perturbation of the Curzon-like structure, as it would be the dominant contribution at short distances from the object. The Weyl multipole expansion is perfectly adapted to analyze this situation. In this case we would fall down into the previous discussion: obtaining (close to) Curzon objects with no trans-Planckian curvatures requires very massive objects with very large cores.
Figure 10: The figure represents, in Weyl coordinates, how generic static and axisymmetric object can become more and more compact approaching the Curzon family of solutions i.e. a combination of eventually singular multipoles. Here we portrait a monopole plus some small deviations (that is, some unspecified Weyl multipoles).
Figure 11: The figure represents, in Weyl coordinates, how generic static and axisymmetric object can become more and more compact its surface narrowing into an infinitely thin rod as it approaches the unique Schwarzschild solution.
@@
Figure 12: The figure represents, in Weyl coordinates (only $(r,z)$ diagram), the difference between a spherically-symmetric ultracompact star, in the left; and a distorsion with respect to it in the right. We have included for reference the singular homogeneous rod source associated with the pure Schwarzschild solution. A spherically-symmetric ultracompact star is such that its surface is located precisely on one of the equipotential surfaces of the Schwarzschild solution close to the rod. A non-spherically symmetric object has distorted equipotential surfaces with respect to the Schwarzschild ones.
In this subsection we first consider a spherically symmetric ultracompact stellar-like object and then introduce in its source inhomogeneous variations in the $z$ direction (see figure 12). Distortions from a spherically symmetric ultracompact object can be obtained by analyzing a potential function $U$ which has the following leading order behaviour $f(z)\log r$ as $r\rightarrow 0$ for $z\in(-M,M)$. The $z$ dependence encodes in the deviations from spherical symmetry. We can introduce an inhomogeneity parameter as $p(z):=f(z)-1$ (also a measurement of the deviation from spherical symmetry), which needs to decrease as we make $\epsilon$ smaller in order to avoid the generation of some trans-Planckian curvatures. At leading order we have that
@@
Here $r_{e}=r_{e}(\epsilon,z)$ is the location of the innermost equipotential surface outside the object at a given $z$. Again we see that for $\epsilon\ll 1$ we need a large ratio mass over size. Now, the leading order behaviour of the Kretschmann for this kind of configurations is
@@
Using (59) the limit on the curvature leads to the condition
@@
For $f(z)=1$ the value of $r_{e}$ is not constrained, but $f(z)=1$ means that the configuration is spherically symmetric. Translated into physical terms, the absence of constraint means that a strictly spherically symmetric object can be as compact as desired without incurring in any large curvature, as long as it remains macroscopic $M\gg\ell_{P}^{-1}$. Now the bigger the departure of $f(z)$ from unity, the larger would have to be the ultracompact object if one does not want to generate trans-Planckian curvatures.
There is alternative way to state the previous result. Let us consider that the mass of the object is fixed and also that $\epsilon$ is fixed. Then, because of (59) the value of $r_{e}$ is also fixed. Now condition (62) must be read from right to left
@@
that is, the non-sphericities of the ultracompact objects must be smaller than the quantity $r_{e}/\ell_{P}$, which in turn is proportional to $\epsilon$ (to some positive power). Then, as the object becomes more and more compact its degree of non-sphericity has to die off or trans-Planckian curvatures will be generated.
We think that the inhomogeneities in the rod that we have just discussed are equivalent to deviations from spherical symmetry. From the point of view of a Geroch-Hansen mass multipolar expansion Geroch (1970); Hansen (1974) , the spherically symmetric configuration corresponds to a pure monopole. However, we have found it difficult to establish the explicit mathematical connection between the inhomogeneous parameter $f(z)$ and the Geroch-Hansen mass multipoles. Far enough from the matter sources the Geroch-Hansen multipoles are uniquely determined in terms of the Weyl multipoles $M^{(\ell)}$ that we have been using before, although there is not an explicit closed expression in terms of them Fodor et al. (1989) and the mapping is non-linear. Generically, the expansion is such that the Geroch-Hansen multipole of order $\ell$ is always composed of a sum of multipoles of the form
$M_{GH}^{(\ell)}=\sum_{n=1}^{\ell}\sum_{k_{1}...k_{n}}c_{k_{1}...k_{n}}M^{(k_{1})}\times\cdots\times M^{(k_{n})},\qquad\sum_{i=1}^{n}k_{i}=\ell$ (64)
where the second equality is a constraint over the sum $\sum_{k_{1}...k_{n}}$ and $c_{k_{1}...k_{n}}$ are coefficients that can be found at any finite order through the algorithm presented in Fodor et al. (1989). However, close to the sources both multipoles represent rather different entities. For instance, it is evident that a pure Geroch-Hansen monopole (i.e., the Schwarzschild metric) cannot be described as a sum of Weyl multipoles everywhere.
@@
There are two previous results in this direction that are worth mentioning and comparing with our findings. In the analysis presented in Barceló et al. (2019), the static case is studied without additional assumptions. It is shown that, under the same conditions imposed here, the Geroch-Hansen multipole structure of the configuration parametrically approaches that of the Schwarzschild solution as the redshift $\epsilon$ tends to zero. However, the paper was restricted to a perturbative analysis of the possible deformations around the Schwarzschild solution which have a smooth limit to it in the $\epsilon\to 0$ limit. This leaves the set of nonperturbative solutions that do not have such a smooth limit outside of the analysis, which in the case under consideration here are easily analyzed since they constitute the whole Curzon sector. In the static and axisymmetric case considered here, we have been able to completely constrain the structure of different ultracompact metrics including deformations of spherically symmetric ultracompact spheroids, which are the ones analyzed in those references.
@@
Something similar happens with the analysis in Raposo et al. (2019). They study linear perturbations on top of a patch of the Schwarzschild metric. Specifically, they consider the Schwarzschild solution with mass $M$ from a surface of proper radius slightly larger than the would-be event horizon, i.e., the region between $r=2M(1+\epsilon)$ and $r\to\infty$. It is shown that the boundedness of the curvature invariants leads to a parametric suppression of the Geroch-Hansen multipole deviations from the Schwarzschild solution. They find that all Geroch-Hansen mass multipoles vanish polynomially in $\epsilon$ if they are “non-spin induced”, meaning that they vanish as the first current multipole of the leading-order perturbation goes to zero. Since we restrict ourselves to the static case, these “non-spin induced” multipoles correspond to the same notion we have analyzed in here. Thus, in the intersecting domain of validity of these analyses, the conclusions are equivalent.
VI Discussion and conclusions
@@
In this work we have analyzed in an exhaustive manner the uniqueness characteristics of black holes under the restricted assumptions of staticity, axisymmetry and asymptotic flatness. One the one hand, we have presented results scattered in the literature in a systematic and novel manner. On the other hand, we have extended these results to black holes immersed in an external gravitational environment and to horizonless ultracompact objects.
@@
The axisymmetric case allows one to show that the only vacuum and asymptotically flat black hole solution with a regular horizon is Schwarzschild spacetime. When compared with Israel theorem Israel (1967) for static geometries, we clearly see that the axisymmetry restriction allows to take rid of two premises of the theorem: the spheroidal topology of the equipotential surfaces condition, and the finite horizon area condition.
@@
Then, we have discussed the difference between deforming a black hole from the inside or from the outside. We have clearly seen that, while it is not possible to deform a black hole horizon from the inside without making it singular, it can indeed be deformed from the outside. These latter deformations are caused by the gravitational fields of external environments. To the best of our knowledge, the only prior research in this direction is by Gürlebeck Gürlebeck (2014, 2015), who demonstrated that the Weyl multipoles for a static and axisymmetric configuration can be expressed as a sum of surface integrals surrounding both the matter and black hole regions. In their interpretation, since the contribution to the multipolar structure from the black hole—i.e., the “hair”—is captured by a single parameter, they conclude that this constitutes a version of the no-hair theorem for black holes in external gravitational fields. However, we argue that this statement can be expressed in a manner closer to more traditional uniqueness results. Our results demonstrate that in a static and axisymmetric scenario, there exists at most a one-parameter family of horizon shapes that can coexist with an external gravitational field (or a given matter configuration). This parameter represents an overall size for the black hole (i.e. it represent its mass). For the environment to be consistent we have shown that it necessary that it exhibits a gravitational equilibrium condition, as worked out in detail in Section III. Therefore, we can say that from a pure geometric (or kinematic) point of view one can have any shape for the black hole horizon. All the uniqueness results come from the dynamical requirements imposed by the vacuum Einstein equations, along with the boundary conditions of the entire gravitating system, one of which is asymptotic flatness.
@@
At this point, we find it necessary to clarify a common misconception about the deformability of black holes, leveraging the results of Gürlebeck. Love numbers, well established in Newtonian gravity and with some working definitions in general relativistic contexts Binnington and Poisson (2009); Poisson (2021), are known to vanish for static black holes Damour and Nagar (2009); Damour and Lecian (2009); Kol and Smolkin (2012); Charalambous et al. (2021). For a static object with a specific multipolar structure deformed by an external gravitational field, Love numbers quantify the extent to which the object’s contribution to the multipolar structure deviates in response to the strength of the external field. As we have shown, this interpretation holds, at least in the static and axisymmetric case under consideration here. However, the vanishing of Love numbers must not be confused with a non-deformability of the cross-sectional geometry of event horizons. Indeed, as the solutions presented here demonstrates, it is evident that the intrinsic geometry of the horizon cross section is deformed due to the presence of external matter. Furthermore, this suggests that one might be able to define some different non-vanishing “Love numbers” to specify how the contribution of the external matter affects the multipolar shape of the deformed horizons.
@@
Finally, we have explored how the no-hair results are modified when relaxing the assumption of the presence of a regular horizon, replacing it with a minimum redshift surface at which the Kretschmann scalar is demanded to be smaller than Planckian. In the static and axisymmetric case considered here, we have been able to constraint all the possible ultracompact structures, including deformations of spherically symmetric ultracompact stars, by requiring that the Kretschmann scalar does not reach trans-Planckian values. For deviations from the Schwarzschild metric described by an inhomogeneity parameter in the direction of the symmetry axis, we find that these inhomogeneities (which are equivalent to non-sphericities) have to die off as the structure becomes more and more compact. In the domains of validity where our analyses intersect with thosein Barceló et al. (2019) and Raposo et al. (2019), we find essentially the same results: that deviations from sphericity have to die off with powers of the compactness parameter. In addition, our analysis has been able to disentangle the role played by the Curzon family of metrics, unveiling why some compactness limits do not lead to the Schwarzschild geometry. We believe that this work would help to better understand the physics behind uniqueness theorems and to better appreciate the additional difficulties in the general static and stationary cases.

### v3

@@ V Almost-no-hair results for horizonless ultracompact objects
As a final case, let us examine the most relevant situation from the astrophysical point of view. It corresponds to ultracompact stellar-like objects that can mimic most properties of general relativistic black holes due to having a deep gravitational well, which makes them almost observationally indistinguishable from black holes. Consider first an ultracompact but spherically symmetric object (see the left-hand side of Fig.  10). Its surface, which is a constant redshift surface $\epsilon$, represents an ellipsoid in Weyl coordinates. However, we know that this is an artifact of using Weyl coordinates and that, from the point of view of the spacetime geometry, it constitutes a perfectly spherically symmetric surface. To prevent trans-Planckian curvatures in this spherically symmetric setup, it suffices for the object to be macroscopic, meaning $M\gg M_{P}$ (which, in our units $G=c=1$, corresponds to $M\gg l_{P}$). The compactness parameter $\epsilon$ becomes irrelevant when we consider only spherically symmetric configurations, as these can be made arbitrarily compact without incurring excessively high curvature values.
@@
Figure 10: The figure, presented in Weyl coordinates [using only the $(r,z)$ plane], illustrates the contrast between a spherically symmetric ultracompact star (left) and a distorted configuration (right). For reference, we have included the homogeneous rod source representing the pure Schwarzschild solution. A spherically symmetric ultracompact star is one whose surface (blue line) precisely aligns with an equipotential surface of the Schwarzschild solution close to the rod (coinciding red line). By contrast, a non-spherically symmetric object exhibits equipotential surfaces that are distorted relative to the Schwarzschild case, and its surface does not necessarily coincide with any equipotential surface. In the right panel, the red line represents the innermost equipotential surface external to the material body.
Now, let us consider deviations from spherical symmetry. We define an object as ultracompact if, outside its matter content, the innermost equipotential surface takes on a very small value, $\epsilon\ll 1$. This surface is distorted compared to the equipotential surfaces of spherically symmetric objects (see the right panel in Fig. 10). Furthermore, in general, the object’s surface does not necessarily coincide with an equipotential surface. In GR, and in the static case, deviations from sphericity are effectively described using the Geroch multipoles Geroch (1970) (in stationary situations these multipoles generalize to the so-called Geroch-Hansen multipoles Hansen (1974)). The set of multipoles consistent with axisymmetry form a series characterized by a single number $\ell$. The strength of each multipole is characterized by a real number $M_{\rm G}^{(\ell)}$, with the monopolar contribution representing the mass of the object, $M_{\rm G}^{(0)}=M$.
It is interesting to point out a subtlety in the relation between Weyl and Geroch multipoles. Far from a matter source, an expansion in Weyl multipoles can always be mapped to an expansion in terms of Geroch multipoles. In particular, a Geroch multipole of order $\ell$ can be decomposed as a sum of Weyl multipoles [defined in Eq. (36)]:
$M_{\rm G}^{(\ell)}=\sum_{n=1}^{\ell}\sum_{k_{1}...k_{n}}c_{k_{1}...k_{n}}M^{(k_{1})}\times\cdots\times M^{(k_{n})};\qquad\sum_{i=1}^{n}k_{i}=\ell,$ (57)
where the second equality is a constraint over the sum $\sum_{k_{1}...k_{n}}$, and $c_{k_{1}...k_{n}}$ are coefficients, for which there is not an explicit closed expression, but which can be found at any finite order through the algorithm presented in Fodor et al. (1989). However, the entire field structure associated with a Geroch multipole, i.e., the full exterior region of the Schwarzschild metric and not only the region far from the event horizon, cannot be described everywhere outside the event horizon as a Weyl multipolar expansion. Specifically, the function $U_{S}$ cannot be expressed as a series of Weyl multipoles centered at a single point. In fact, this shows that the Weyl multipolar expansion is not well-suited for describing spherically symmetric configurations
It is evident that different ultracompact material distributions can produce the same external multipolar expansion. In other words, a given gravitational field generated by a matter distribution (noting that there is no horizon, ensuring that it is causally connected with the exterior region) can arise from infinitely many different internal matter configurations. However, for isolated objects and in the singular $\epsilon\to 0$ limit, the distortions (which actually become infinite on the $\epsilon=0$ surface) from the pure Schwarzschild form can only come about essentially in two essential forms. Ultracompact objects either tend to an inhomogeneous rod (see the pictorial representation on the left of Fig. 11) or they acquire some additional Weyl multipolar contributions localized at isolated points of the rod (see the illustrative drawing on the right of Fig. 11). It all depends on whether making $\epsilon$ smaller unveils a $f(z)\log(r)$ behavior of the redshift function or the $1/r^{\ell+1}$ behavior associated with a Weyl multipole contribution. In the latter case, the geometry is dominated by the Weyl multipoles as $\epsilon\to 0$ so it would be more appropriate to talk about distorted Curzon objects than to talk about distorted Schwarzschild black holes. The curvature condition will impose on these objects the constraints that we already discussed for ultracompact Curzon-like objects. Hence, we focus now on ultracompact objects whose leading order behavior for $U$ is $f(z)\log r$ as $r\rightarrow 0$ for $z\in(-M,M)$.
Figure 11: The red lines in the figure represent in Weyl coordinates the innermost equipotential surfaces outside the body for different ultracompact objects,. The closer the surface to the axis of symmetry the bigger the compactness of the object. The figure on the left represents a series of more and more compact objects tending to a limit described by an inhomogeneous rod, with a linear density function $\lambda(z)$. The figure on the right represents instead a series of ultracompact configurations which, apart from an inhomogeneous rod, unveil a singular Weyl monopolar contribution. Essentially, these are the two types of tendencies one could find when analyzing the geometry of ultracompact objects with higher and higher compactness.
As we are going to see, $f(z)$ needs to become closer to $1$ everywhere in $z\in(-M,M)$ as we make $\epsilon$ smaller in order to avoid the generation of trans-Planckian curvatures. If we think of a distortion with respect to an ultracompact spherically-symmetric object of mass $M$, we can impose additional constraints on $\lambda(z)$. In particular, we can impose that: i) the total mass of the system is fixed, and ii) that its center of mass is located at the origin of coordinates $(r,z)=(0,0)$. This is equivalent to:
$\int_{-M}^{+M}dz\,[f(z)-1]=0;~~~~~\int_{-M}^{+M}dz\,[f(z)-1]z=0.$ (58)
For an ultracompact stellar-like object we have, at leading order in the limit $\epsilon\to 0$,
@@
Here, $r_{e}=r_{e}(\epsilon,z)$ is the location of the innermost equipotential surface outside the object at a given $z$. We see again that for $\epsilon\ll 1$ we need a large mass-over-size ratio. Now, the leading order behavior of the Kretschmann scalar for this kind of configurations is
@@
Using Eq. (59), the curvature bound leads to the condition
@@
For $f(z)=1$ the value of $r_{e}$ is not constrained, which is precisely the case in which the configuration is spherically symmetric. Hence, a strictly spherically symmetric object can be as compact as desired without incurring in Planckian curvatures, as long as it remains macroscopic ($M\gg\ell_{P}$). On the other hand, when $f(z)\neq 1$, the bigger the departure of $f(z)$ from unity, the larger would have to be the ultracompact object if one does not want to generate trans-Planckian curvatures. Notice that Eq. (58), which fixes the monopole, automatically excludes the case in which $f(z)=0$ everywhere. Actually, $f(z)$ can become zero at some points, although for instance, if it is a simple zero, the subleading term in Eq. (43) would still impose constraints on $f^{\prime}(z)$. This remains true for some values of $z$ even in the case in which $f(z)=0$ is zero on an interval $I\subset(-M,M)$, which physically would represent two disconnected objects separated by a vacuum region in between.
There is an alternative way to state the previous result. Let us consider that both the mass of the object and $\epsilon$ are fixed. Then, due to Eq. (59) the value of $r_{e}$ is also fixed. Now Eq. (62) must be read from right to left:
@@
That is, the non-sphericities of the ultracompact object must be smaller than the quantity $r_{e}/\ell_{P}$, which in turn is proportional to (some positive power of) $\epsilon$. Then, as the object becomes more and more compact, its degree of non-sphericity has to die off or trans-Planckian curvatures will be generated.
There needs to exist a one-to-one correspondence between the functions $f(z)$ and the Geroch multipoles $M_{\rm G}^{(\ell)}$, with $\ell>1$, of the ultracompact object. The specific functional relation can be however quite convoluted and we do not aim to present it here. We will just sketch for the sake of completeness how it could be established, and leave it for future work.
First of all, we notice that the Weyl form of the Schwarzschild metric in Eq. (38) can be transformed into the standard Schwarzschild coordinates $(u,\theta)$ by performing the transformation
$r=\sqrt{u(u-2M)}\sin\theta;~~~~~z=(u-M)\cos\theta.$ (64)
Applying this transformation blindly to a generic Weyl metric (28) we obtain
$ds^{2}=$ $-e^{2U}dt^{2}$
$+e^{-2U}\bigg{[}e^{2V}\left((1+\frac{M^{2}\sin^{2}\theta}{u^{2}-2Mu})du^{2}+\left(u^{2}-2Mu+M^{2}\sin^{2}\theta\right)d\theta^{2}\right)$
$+u(u-2M)\sin^{2}\theta d\varphi^{2}\bigg{]},$ (65)
where the functions $U$ and $V$ need to be expressed in terms of $u$ and $\theta$. The potential $U$ associated with an inhomogeneous rod of density $\lambda(z)$ is given by
$U(r,z)=-\int_{-M}^{M}dz^{\prime}\frac{\lambda(z^{\prime})}{\sqrt{r^{2}+(z^{\prime}-z)^{2}}}.$ (66)
Let us recall that, for a fixed $z\in(-M,M)$ in the limit $r\to 0$, we have
$U(r,z)\simeq 2\lambda(z)\ln r=f(z)\ln r.$ (67)
As discussed at the beginning of the article, from $U$ one can directly obtain $V$ through quadratures (although in general they cannot be written in closed form) and then plug them into the line element in Eq. (65). Because of the previous conditions on the center of mass, we can be sure that the Schwarzschild coordinates that we are using are Thorne’s “Asymptotically Cartesian and Mass-Centered” (ACMC) coordinates Thorne (1980). Then, from the form of the metric one could extract Thorne’s multipoles, and in turn Geroch’s multipoles Gürsel (1983). Deviations from the Schwarzschild monopolar contribution that preserve the center of mass are described by the $\ell>1$ Geroch multipole decomposition. This establishes the connection between the inhomogeneity function $\lambda(z)$ and Geroch multipoles. It is then clear that Geroch multipoles have to die off as compactness becomes very small to avoid the generation of trans-Planckian curvatures, as the function $f(z)$ tends to $1$.
@@
Let us now mention and briefly compare with our findings two previous results in the same direction. In the analysis presented in Barceló et al. (2019), the static case is studied without additional assumptions. There it is shown that, under the same conditions imposed here, the Geroch multipole structure of the configuration parametrically approaches that of the Schwarzschild solution as the redshift $\epsilon$ tends to zero. The paper was restricted to a perturbative analysis of the possible deformations around the Schwarzschild solution that have a smooth limit to it in the $\epsilon\to 0$ limit. This leaves the set of nonperturbative solutions that do not have such a smooth limit outside the analysis, which in the case under consideration here are easily analyzed since they constitute the whole Curzon sector. In the static and axisymmetric case considered here, we have been able to completely constrain the structure of different ultracompact metrics including deformations of spherically-symmetric ultracompact spheroids, which are the ones analyzed in this reference.
@@
Something similar happens with the analysis in Raposo et al. (2019). They study linear perturbations on top of a patch of the Schwarzschild metric. Specifically, they consider the Schwarzschild solution with mass $M$ from a surface of proper radius slightly larger than the would-be event horizon, i.e., the region between $r=2M(1+\epsilon)$ and $r\to\infty$. There it is shown that the boundedness of curvature invariants leads to a parametric suppression of the Geroch multipole deviations from the Schwarzschild solution. They find that all Geroch-Hansen mass multipoles vanish polynomially in $\epsilon$ if they are “non-spin induced”, meaning that they vanish as the first current multipole of the leading-order perturbation goes to zero. Since we restrict ourselves to the static case, these “non-spin induced” multipoles correspond to the same notion we have analyzed in here. Thus, in the intersecting domain of validity of these analyses, their conclusions and ours are equivalent.
To conclude this section, it is important to highlight that, for an ultracompact object with a small but fixed $\epsilon,$ it is not strictly possible to determine whether the distortions from strict sphericity originate from within the object or from external influences, as both sources would lead to finite distortions. To distinguish between these possibilities, one would need to probe the external gravitational background and assess whether the object exhibits distortions beyond those naturally caused by external gravitational effects. What we have demonstrated here is that any distortion originating from within the object must be very small—the smaller the more compact the body is.
VI Conclusions and further remarks
@@
In this work we have analyzed in an exhaustive manner the uniqueness characteristics of black holes under the assumptions of staticity, axisymmetry, and asymptotic flatness. One the one hand, we have presented results scattered in the literature in a systematic and novel manner. On the other hand, we have extended these results to black holes immersed in an external gravitational environment and to horizonless ultracompact objects.
@@
The axisymmetric case allows one to show that the only vacuum and asymptotically flat black hole solution with a regular horizon is Schwarzschild spacetime. When compared with Israel theorem Israel (1967) for static geometries, we clearly see that the axisymmetry restriction allows to get rid of two premises of the theorem: the spheroidal topology of the equipotential surfaces condition, and the finite horizon area condition.
@@
Then, we have discussed the difference between deforming a black hole from the inside or from the outside. We have clearly seen that, while it is not possible to deform a black hole horizon from the inside without making it singular, it can indeed be deformed from the outside. These latter deformations are caused by the gravitational fields of external environments. To the best of our knowledge, the only prior research in this direction is by Gürlebeck Gürlebeck (2014, 2015), who demonstrated that the Weyl multipoles for a static and axisymmetric configuration can be expressed as a sum of surface integrals surrounding both the matter and black hole regions. In their interpretation, since the contribution to the multipolar structure from the black hole—i.e., the “hair”—is captured by a single parameter, they conclude that this constitutes a version of the no-hair theorem for black holes in external gravitational fields. However, we argue that this statement can be expressed in a manner closer to more traditional uniqueness results. Our results demonstrate that in a static and axisymmetric scenario, there exists at most a one-parameter family of horizon shapes that can coexist with an external gravitational field (or a given matter configuration). This parameter represents an overall size for the black hole (i.e. it represents its mass). For the environment to be consistent we have shown that it is necessary that it exhibits a gravitational equilibrium condition, as worked out in detail in Section III. Therefore, we can say that from a pure geometric (or kinematic) point of view one can have any shape for the black hole horizon. All the uniqueness results come from the dynamical requirements imposed by the vacuum Einstein equations, along with the boundary conditions of the entire gravitating system, one of which is asymptotic flatness being one of them.
@@
At this point, we find it necessary to clarify a common misconception about the deformability of black holes, leveraging the results of Gürlebeck. Love numbers, well established in Newtonian gravity and with some working definitions in general relativistic contexts Binnington and Poisson (2009); Poisson (2021), are known to vanish for static black holes Damour and Nagar (2009); Damour and Lecian (2009); Kol and Smolkin (2012); Charalambous et al. (2021). For a static object with a specific multipolar structure deformed by an external gravitational field, Love numbers quantify the extent to which the object’s contribution to the multipolar structure deviates in response to the strength of the external field. As we have shown, this interpretation holds, at least in the static and axisymmetric case under consideration here. However, the vanishing of Love numbers must not be confused with a non-deformability of the cross-sectional geometry of event horizons. Indeed, as the solutions presented here demonstrate, it is evident that the intrinsic geometry of the horizon cross section is deformed due to the presence of external matter. Furthermore, this suggests that one might be able to define some different non-vanishing “Love numbers” to specify how the contribution of the external matter affects the multipolar shape of the deformed horizons.
Our explanation of why black holes cannot be distorted from the inside is somewhat technical, so here we offer a more intuitive, physical interpretation. A lump of matter can resist its own gravitational pull by generating internal positive pressures that counteract it. When an object is not very compact, this internal pressure largely determines its shape. For example, objects such as asteroids with very low compactness can take on very irregular, non-spherical shapes. However, as an object’s compactness increases, gravity becomes a dominant factor in shaping it, and non-rotating objects tend to become more spherical. This spherical shape minimizes the object’s gravitational self-energy. When the compactness of an object surpasses the black-hole threshold (i.e., when an event horizon forms), any structure capable of resisting gravitational collapse would require violating energy conditions—either by creating negative pressures or negative energy densities. For instance, one could construct a spherically symmetric black holes that only differ from the Schwarzschild black hole in an inner core located in the black hole region Frolov et al. (1989). In fact, any non-spherically symmetric regular black hole would require external matter to sustain the deviation from spherical symmetry. This can be understood as follows: consider a geometry with a non-vanishing $T_{\mu\nu}$ in an inner core that is not spherically symmetric but depends continuously on a parameter representing its compactness. As the configuration becomes progressively more compact, there will be a point where a horizon forms. However, our analysis indicates that such a horizon would inevitably leave some matter outside. If the horizon were to enclose all the matter content, it would necessarily result in a spherically symmetric horizon.
@@
Finally, in this paper we have explored how the no-hair results are modified when relaxing the assumption of the presence of a regular horizon, replacing it with a minimum redshift surface at which the Kretschmann scalar is demanded to be smaller than Planckian. In the static and axisymmetric case considered here, we have been able to constrain all the possible ultracompact structures, including deformations of spherically symmetric ultracompact stars by requiring that the Kretschmann scalar does not reach trans-Planckian values. For deviations from the Schwarzschild metric described by an inhomogeneity function in the direction of the symmetry axis, we find that these inhomogeneities (which are equivalent to non-sphericities) have to die off as the structure becomes more and more compact. In the domains of validity where our analyses intersect with those in Barceló et al. (2019) and Raposo et al. (2019), we find essentially the same results: that deviations from sphericity have to die off with powers of the compactness parameter. In addition, our analysis has been able to disentangle the role played by the Curzon family of metrics, unveiling why some compactness limits do not lead to the Schwarzschild geometry. We believe that this work would help to better understand the physics behind uniqueness theorems and to better appreciate the additional difficulties in the general static and stationary cases.